# On Neural Networks as Infinite Tree-Structured Probabilistic Graphical Models

### v3

@@ 1 Introduction
Deep neural networks (DNNs), including large language models, offer state-of-the-art prediction performance, but they are difficult to interpret due to their complex multilayer structure, large number of latent variables, and the presence of nonlinear activation functions (Buhrmester et al., 2021). To gain a precise statistical interpretation for DNNs, much progress has been made in linking them to probabilistic graphical models (PGMs). Variational autoencoders (VAEs) (Kingma and Welling, 2014) are an early example; more recent examples relate recurrent neural networks (RNNs) with hidden Markov models (HMMs) (Choe et al., 2017) and convolutional neural networks (CNNs) with Gaussian processes (GPs) (Garriga-Alonso et al., 2018). When such a connection is possible, benefits include:
@@
  • Clear statistical semantics for a trained DNN model beyond providing the conditional distribution over output variables given input variables. Instead, PGMs provide a joint distribution over all variables including latent variables.
  • Ability to make inferences about how evidence of some nodes influences probabilities at others, including how later nodes influence earlier ones, as in Bayes nets or Markov nets.
  • Ability to understand weight initializations of DNNs as representing prior distributions and trained DNNs as representing posterior distributions, or ensembles of models.
  • Proposal of new algorithms by importing algorithmic approaches from PGMs into DNNs.
@@
In this paper, we establish a correspondence between DNNs and PGMs. Given an arbitrary DNN, we first construct an infinite-width tree-structured PGM. We then demonstrate that during training, the DNN executes approximations of precise inference in the PGM during the forward propagation. We prove our result in the case of sigmoid activations and then indicate how the proof can be expanded to other activation functions, provided that some form of normalization is employed. These findings provide immediate benefits such as those listed above.
@@
This work stands apart from most theoretical analyses of DNNs, which typically view DNNs purely as function approximators and prove theorems about the quality of function approximation. Here we instead show that DNNs may be viewed as statistical models, specifically PGMs. This work is also different from the field of Bayesian neural networks, where the goal is to seek and model a probability distribution over neural network parameters. In our work, the neural network itself defines a joint probability distribution over its variables (nodes). Our work therefore is synergistic with Bayesian neural networks but more closely related to older work to learn stochastic neural networks via expectation maximization (EM) (Amari, 1995) or approximate EM (Song et al., 2016).
@@
Although the approach is different, our motivation is similar to that of Dutordoir et al. (2021) and Sun et al. (2020) in their work to link DNNs to deep Gaussian processes (GPs) (Damianou and Lawrence, 2013). By identifying the forward pass of a DNN with the mean of a deep GP layer, they aim to augment DNNs with advantages of GPs, notably the ability to quantify uncertainty over both output and latent nodes. What distinguishes our work is that we make the DNN-PGM approximation explicit and include all sigmoid DNNs, not just unsupervised belief networks or other specific cases.
@@ 2 Background: Comparison to Bayesian Networks and Markov Networks
Syntactically a Bayesian network (BN) is a directed acyclic graph, like a neural network, whose nodes are random variables. Semantically, a BN represents a full joint probability distribution over its variables as $P(\vec{v})=\prod_{i}P(v_{i}|pa(v_{i}))$, where $\vec{v}$ is a complete setting of the variables, and $pa(v_{i})$ denotes the parents of variable $v_{i}$. If the conditional probability distributions (CPDs) $P(v_{i}|pa(v_{i}))$ are all logistic regression models, we refer to the network as a sigmoid BN.
@@
A DNN of any architecture is, like a Bayesian network, a directed acyclic graph. A sigmoid activation can be understood as a logistic model, thus giving a conditional probability distribution for a binary variable given its parents. Thus, there is a natural interpretation of a DNN with sigmoid activations as a Bayesian network (e.g., Bayesian belief network).***Note that when viewing the DNN as a function approximator, these conditional probability distributions are not explicitly defined. As reviewed in theorem 1, this Bayes net in turn is equivalent to (represents the same probability distribution) as a Markov network where every edge of weight $w$ from variable $A$ to variable $B$ has a potential of the following form:
@@
Theorem 1.
We don’t claim theorem 1 is new, but we provide a proof in Appendix A because it captures several components of common knowledge to which we couldn’t find a single reference.
@@ 3 The Construction of Tree-structured PGMs
Although both a binary pairwise Markov network (MN) and a Bayesian network (BN) share the same sigmoid functional structure as a DNN with sigmoid activations, it can be shown that the DNN does not in general define the same probability for the output variables given the input variables: forward propagation in the DNN is very fast but yields a different result than VE in the MN or BN, which can be much slower because the inference task is NP-complete. Therefore, if we take the distribution $\cal{D}$ defined by the BN or MN to be the correct meaning of the DNN, the DNN must be using an approximation $\cal{D^{\prime}}$ to $\cal{D}$. Procedurally, the approximation can be shown to be exactly the following: the DNN repeatedly treats the expectation of a variable $V$, given the values of $V$’s parents, as if it were the actual value of $V$. Thus previously binary variables in the Bayesian network view and binary features in the Markov network view become continuous. While this procedural characterization of the approximation of $\cal{D^{\prime}}$ to $\cal{D}$ is precise, we prefer in the PGM literature to characterize approximate distributions such as $\cal{D^{\prime}}$ with an alternative PGM that precisely corresponds to $\cal{D^{\prime}}$; for example, in some variational methods we may remove edges from a PGM to obtain a simpler PGM in which inference is more efficient. Treewidth-1 (tree-structured or forest-structured) PGMs are among the most desirable because in those exact inference by VE or other algorithms becomes efficient. We seek to so characterize the DNN approximation here.
@@
  3. 3.
    Consider then the children of these nodes. Again, for each outgoing edge, make a copy of this node that is only connected to one child with that edge. In this step, for each copied node, we then also copy the entire subgraph formed by all ancestor nodes of the current node. Note that while weights across copies are tied, the values of the copies of any node are not tied. However, since we also copy the subtree of all input and intermediary hidden nodes relevant in the calculation of this node for each copy, the probability of any of these copied nodes being true remains the same across copies.
  4. 4.
    We repeat this process until we have separate trees for each output node in the original deep neural network graph.
This process ultimately creates a graph whose undirected structure is a tree or forest. In the directed structure, trees converge at the output nodes. The probability of any copy of a latent node given the observed input is the same across all the copies, but when sampling, their values may not be the same.
@@
The preceding step alone is still not sufficient to accurately express the deep neural network as a PGM. Recall that in the Markov network view, we have seen that the neural network makes a mean-field approximation where it uses the expected value of a node in place of its actual value. The following additional step in the construction yields this same behavior. This next step of the construction creates $L$ copies of every non-output node in the network while also copying the entire subtrees of each of these nodes, as was done in step 1. The weight of a copied edges is then set to its original value divided by $L$. As $L$ approaches infinity, we show that the gradient in this PGM construction matches the gradient in the neural network exactly.
@@
This second step in the construction can be thought of intuitively by considering the behavior of sampling in the Bayesian network view. Since we make $L$ copies of each node while also copying the subgraph of its ancestors, these copied nodes all share the same probabilities. As $L$ grows large, even if we sampled every copied node only once, we would expect the average value across these $L$ copies to match the probability of an individual copied node being true. Given that we set the new weights between these copies and their parents as the original weights divided by $L$, the sum of products (new weights times parent values) yields the average parent value multiplied by the original weight. As $L$ goes to infinity, we remove sampling bias and the result exactly matches the value of the sigmoid activation function of the neural network, where this expectation in the PGM view is passed repeatedly to the subsequent neurons. The formal proof of this result, based on variable elimination, is found below. There, we show the following:
Theorem 2.
@@
In the PGM construction, as $L\to\infty$, $P(H=1|\vec{x})\to\sigma(\sum_{j=1}^{M}w_{j}g_{j}+\sum_{i}^{N}\theta_{i}\sigma(p_{i}))$, for an arbitrary latent node $H$ in the DNN that has observed parents $g_{1},...,g_{M}$ and latent parents $h_{1},...,h_{N}$ that are true with probabilities $\sigma(p_{1}),...,\sigma(p_{N})$. $w_{1},...,w_{M}$ and $\theta_{1},...,\theta_{N}$ are the weights on edges between these nodes and $H$.
In order to prove that as $L$ goes to infinity, this PGM construction does indeed match the neural network’s forward propagation, we consider an arbitrary latent node $H$ with $N$ unobserved parents $h_{1},...,h_{N}$, $M$ observed parents $g_{1},...,g_{M}$, and a single child node $C$. The edges between these parents and $H$ then each have a weight $\theta_{i}$, $1\leq i\leq N$, for the unobserved nodes, and $w_{j}$, $1\leq j\leq M$, for the observed nodes, and $\gamma$, for the child node. The network as a whole has observed evidence $\vec{x}$. For the rest of this problem we use a Markov network view of the neural network. The potentials for these nodes in the network are as follows:
$h_{i}$$\neg h_{i}$
$e^{p_{i}}$ 1
Since $g_{j}$ are observed, their values are found in $\vec{x}$. The potential for the ’child’ node can be represented as:
$C$$\neg C$
$c$ 1
where $c$ is a positive real number that can be arbitrarily large.
$H$$\neg H$
$h_{i}$ $e^{\theta_{i}}$ 1
$\neg h_{i}$ $1$ 1
$H$$\neg H$
$g_{j}$ $e^{w_{j}}$ 1
$\neg g_{j}$ $1$ 1
$H$$\neg H$
$C$ $e^{\gamma}$ 1
$\neg C$ $1$ 1
Suppose, then, using the second step of our construction, we make $L$ copies of all the nodes that were parents of $H$ in the Bayesian network view of the DNN, $h_{1}^{1},...,h_{1}^{L},$ ,...,… , $h_{N}^{1},...,h_{N}^{L}$ and $g_{1}^{1},...,g_{1}^{L},$ ,...,… , $g_{M}^{1},...,g_{M}^{L}$ with weights $\theta_{1}/L,...,\theta_{N}/L$ and $w_{1}/L,...,w_{M}/L$ respectively. Note, as per the construction, additional copies of $C$ are not connected to its parent $H$, however, its weight becomes $\gamma/L$. The potential on $H$ and these copied nodes is then:
$H$$\neg H$
$h_{i}^{k}$ $e^{\theta_{i}/L}$ 1
$\neg h_{i}^{k}$ $1$ 1
$H$$\neg H$
$g_{j}^{k}$ $e^{w_{j}/L}$ 1
$\neg g_{j}^{k}$ $1$ 1
where $1\leq i\leq N$, $1\leq j\leq M$, and $1\leq k\leq L$. The potentials for each of the copied nodes are the same as the nodes they were originally copied from. We then have that,
$P(H,C,h_{1}^{1},...,h_{1}^{L},...,h_{N}^{1},...,h_{N}^{L},g_{1},...,g_{1}^{L},...,g_{M}^{1},...,g_{M}^{L}|\vec{x})$
$=\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times\prod_{j=1}^{M}\prod_{k=1}^{L}e^{(w_{j}/L)H\times g_{j}^{k}}\times\prod_{i=1}^{N}\prod_{k=1}^{L}e^{(\theta_{i}/L)H\times h_{i}^{k}}$
$=\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times e^{H(\frac{\theta_{1}}{L}\sum_{k=1}^{L}h_{1}^{k}+...+\frac{\theta_{N}}{L}\sum_{k=1}^{L}h_{N}^{k})}.$
Summing out an arbitrary, copied latent node, $h_{\alpha}^{\beta}$:
$\sum_{h_{\alpha}^{\beta},\neg h_{\alpha}^{\beta}}P(H,C,h_{1}^{1},...,h_{1}^{L},...,h_{N}^{1},...,h_{N}^{L}|\vec{x})$
$=\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times\sum_{h_{\alpha}^{\beta},\neg h_{\alpha}^{\beta}}\prod_{i=1}^{N}\prod_{k=1}^{L}e^{(\theta_{i}/L)H\times h_{i}^{k}}$
$=\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times$
$\left[e^{p_{\alpha}}e^{(\theta_{\alpha}/L)H}\times\prod_{\begin{subarray}{c}i=1,..,N\\ (i,k)\neq(\alpha,\beta)\end{subarray}}\prod_{k=1,...L}e^{(\theta_{i}/L)H\times h_{i}^{k}}\right.$
$\hskip 56.9055pt\left.+\prod_{\begin{subarray}{c}i=1,..,N\\ (i,k)\neq(\alpha,\beta)\end{subarray}}\prod_{k=1,...L}e^{(\theta_{i}/L)H\times h_{i}^{k}}\right]$
$=\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times(e^{p_{\alpha}}e^{(\theta_{\alpha}/L)H}+1)$
×$\hskip 56.9055pt\times\left[\prod_{\begin{subarray}{c}i=1,..,N\\ (i,k)\neq(\alpha,\beta)\end{subarray}}\prod_{k=1,...L}e^{(\theta_{i}/L)H\times h_{i}^{k}}\right]_{\textstyle.}$
Summing out all $L$ copies of $h_{\alpha}$:
$\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times(e^{p_{\alpha}}e^{(\theta_{\alpha}/L)H}+1)^{L}$
×$\hskip 42.67912pt\times\left[\prod_{\begin{subarray}{c}i=1,..,N\\ i\neq\alpha\end{subarray}}\prod_{k=1,...L}e^{(\theta_{i}/L)H\times h_{i}^{k}}\right]_{\textstyle.}$
Summing out the $L$ copies of each latent parent would then yield:
$\frac{1}{Z}\times e^{(\gamma/L)H\times C}\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)H}+1)^{L}\;,$
We then sum out $C$:
$\frac{1}{Z}\times(ce^{(\gamma/L)H}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times H}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)H}+1)^{L}\;,$
which, in turn, gives us:
$P(H=1|\vec{x})$
$=\left[(ce^{(\gamma/L)\times 1}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times 1}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)\times 1}+1)^{L}\right]\bigg{/}$
$\hskip 11.38109pt\left[(ce^{(\gamma/L)\times 1}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times 1}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)\times 1}+1)^{L}\right.$
$\hskip 17.07164pt\left.+(ce^{(\gamma/L)\times 0}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}\times 0}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)\times 0}+1)^{L}\right]$
$=\frac{(ce^{(\gamma/L)}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)}+1)^{L}}{(ce^{(\gamma/L)}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)}+1)^{L}+(c+1)\prod_{i}^{N}(e^{p_{i}}+1)^{L}}$
$=\left[\frac{1}{1+\frac{(c+1)\prod_{i}^{N}(e^{p_{i}}+1)^{L}}{(ce^{(\gamma/L)}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)}+1)^{L}}}\right]_{\textstyle.}$
We then consider:
$\lim_{L\to\infty}\frac{(c+1)\prod_{i}^{N}(e^{p_{i}}+1)^{L}}{(ce^{(\gamma/L)}+1)\times e^{\sum_{j=1}^{M}w_{j}g_{j}}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)}+1)^{L}}=$
$\lim_{L\to\infty}\exp\left(-\sum_{j=1}^{M}w_{j}g_{j}+\sum_{i=1}^{N}L\times\log(e^{p_{i}}+1)\right.$
$\hskip 51.21495pt\left.-\sum_{i=1}^{N}L\times\log(e^{p_{i}}e^{(\theta_{i}/L)}+1)-\log(ce^{(\gamma/L)}+1)+\log(c+1)\right)_{\textstyle,}$
and the logarithm of this limit is,
$\lim_{L\to\infty}\left[-\sum_{j=1}^{M}w_{j}g_{j}+\sum_{i=1}^{N}L\times\log(e^{p_{i}}+1)\right.$
$\hskip 28.45274pt\left.-\sum_{i=1}^{N}L\times\log(e^{p_{i}}e^{(\theta_{i}/L)}+1)-\log(ce^{(\gamma/L)}+1)+\log(c+1)\right]$
$=-\sum_{j=1}^{M}w_{j}g_{j}\hskip 5.69054pt+\lim_{L\to\infty}\left[-\log(ce^{(\gamma/L)}+1)+\log(c+1)\right]+$
$\hskip 28.45274pt\lim_{L\to\infty}\frac{\sum_{i=1}^{N}\left(\log(e^{p_{i}}+1)-\log(e^{p_{i}}e^{(\theta_{i}/L)}+1)\right)}{1/L}.$
Note that $\lim_{L\to\infty}\left[-\log(ce^{(\gamma/L)}+1)+\log(c+1)\right]=0$ and the second limit in the previous expression clearly has the indeterminate form of $\frac{0}{0}$. Let $G=-\sum_{j=1}^{M}w_{j}g_{j}$ and consider, then the following change of variables, $S=1/L$, and subsequent use of l’Hôspital’s rule.
$G+\lim_{S\to 0^{+}}\frac{\sum_{i=1}^{N}\left(\log(e^{p_{i}}+1)-\log(e^{p_{i}}e^{(\theta_{i}S)}+1)\right)}{S}$
$=G+\lim_{S\to 0^{+}}\frac{\frac{\partial}{\partial S}\sum_{i=1}^{N}\left(\log(e^{p_{i}}+1)-\log(e^{p_{i}}e^{(\theta_{i}S)}+1)\right)}{\frac{\partial}{\partial S}S}$
$=G+\lim_{S\to 0^{+}}\frac{\sum_{i=1}^{N}\frac{-1}{e^{p_{i}}e^{(\theta_{i}S)}+1}\times e^{p_{i}}e^{\theta_{i}S}\times\theta_{i}}{1}$
$=G+\lim_{S\to 0^{+}}\sum_{i=1}^{N}\frac{-e^{p_{i}}e^{\theta_{i}S}\times\theta_{i}}{e^{p_{i}}e^{\theta_{i}S}+1}$
$=G-\sum_{i}^{N}\frac{e^{p_{i}}}{e^{p_{i}}+1}\times\theta_{i}$
$=-\sum_{j=1}^{M}w_{j}g_{j}-\sum_{i}^{N}\sigma(p_{i})\theta_{i}.$
Therefore,
$\lim_{L\to\infty}\frac{\prod_{i}^{N}(e^{p_{i}}+1)^{L}}{e^{\sum_{j=1}^{M}w_{j}g_{j}}\times\prod_{i}^{N}(e^{p_{i}}e^{(\theta_{i}/L)}+1)^{L}}$
$=\exp(-\sum_{j=1}^{M}w_{j}g_{j}-\sum_{i}^{N}\sigma(p_{i})\theta_{i}),$
and,
$\lim_{L\to\infty}P(H=1|\vec{x})$
$=\frac{1}{1+e^{-\sum_{j=1}^{M}w_{j}g_{j}-\sum_{i}^{N}\sigma(p_{i})\theta_{i}}}$
$=\sigma(\sum_{j=1}^{M}w_{j}g_{j}+\sum_{i}^{N}\sigma(p_{i})\theta_{i}).$
This is exactly the result of the deep neural network. Suppose then that $z$ is a hidden node whose parents in the Bayesian network view are all observed. By our PGM construction, we have that the potential for $z$ true is $e^{\sum_{x\in\vec{x}}w_{zx}x}$ where $w_{zx}$ is the weight between nodes $z$ and $x$, and, for $z$ false, 1. This clearly matches the deep neural network’s sigmoid activation in this ‘first layer’. Consider, then, the nodes whose parents in the Bayesian network view are either one of these first layer hidden nodes, or an observed node. By our PGM construction, we have shown that so long as nodes in the previous layer are either observed or have sigmoid conditional probabilities, as is the case here, the conditional probability of any nodes that immediately follow will also have a sigmoid conditional probability. Repeating this argument up to the output nodes gives us that the conditional probability in this PGM construction and the activation values of the DNN match for any layer in the DNN.
Consider then the view of the DNN where each layer is defined such that the values of the activation functions for all neurons in a given layer can be calculated using neurons of the preceding layers. The DNN is structured then such that the first layer depends only on observed evidence and all layers can be calculated sequentially from that starting point. We have already established that nodes with only observed parents have this sigmoid conditional probability. Given the structure of the DNN and Theorem 2, we then have that the corresponding layers in our PGM construction of the DNN can be similarly computed sequentially from that first layer and have conditional probabilities that exactly match the DNN’s activations.

### v5

@@ 1 Introduction
Deep neural networks (DNNs), including large language models, offer state-of-the-art performance on many tasks, but they are difficult to interpret due to their complex structure, large number of latent variables, and the presence of nonlinear activation functions (Buhrmester et al., 2021). To gain a precise statistical interpretation for DNNs, much progress has been made in linking them to probabilistic graphical models (PGMs). Variational autoencoders (VAEs) (Kingma and Welling, 2014) are an early example; more recent examples include probabilistic dependency graphs (Richardson, 2022) as well as work that relates recurrent neural networks (RNNs) with hidden Markov models (HMMs) (Choe et al., 2017) and convolutional neural networks (CNNs) with Gaussian processes (GPs) (Garriga-Alonso et al., 2018). When such a connection is possible, potential benefits include:
@@
  • Clear statistical semantics for a trained DNN model, beyond merely providing the conditional distribution over output variables given input variables. Instead, PGMs provide a joint distribution over all variables including the latent variables.
  • Ability to import any PGM algorithm into DNNs, such as belief propagation or MCMC, for example to reverse any DNN to use output variables as evidence and variables anywhere earlier as query variables.
  • Improved calibration, i.e., improved predictions of probabilities at the output nodes by incorporation of PGM algorithms.
@@
In this paper, we establish such a correspondence between DNNs of any structure and PGMs. Given an arbitrary DNN architecture, we first construct an infinite-width tree-structured PGM. We then demonstrate that during training, the DNN executes approximations of precise inference in the PGM during the forward propagation step. We prove our result exactly in the case of sigmoid activations. We indicate how it can be extended to ReLU activations by building on a prior result of Nair and Hinton (2010). Because the PGM in our result is a Markov network, the construction can extend even further, to all nonnegative activation functions provided that proper normalization is employed. We argue that modified variants of layer normalization and batch normalization could be viewed as approximations to proper MN normalization in this context, although formal analyses of such approximations are left for future work. Finally, in the context of sigmoid activations we empirically evaluate how the second and third benefits listed above follow from the result, as motivated and summarized now in the next paragraph.
A neural network of any architecture with only sigmoid activations satisfies the definition of a Bayesian network over binary variables—it is a directed acyclic graph with a conditional probability distribution at each node, conditional on the values of the node’s parents—and thus is sometimes called a Bayesian belief network (BBN). Nevertheless, it can be shown that the standard gradient used in neural network training (whether using cross-entropy or other common error functions) is inconsistent with the BBN semantics, that is, with the probability distribution defined by this BBN. On the other hand, Gibbs sampling is a training method consistent with the BBN semantics, and hence effective for calibration and for reversing any neural network, but it is woefully inefficient for training. Hamiltonian Monte Carlo (HMC) is more efficient than Gibbs and better fits BBN semantics than SGD. We demonstrate empirically that after training a network quickly using SGD, calibration can be improved by fine-tuning using HMC. The specific HMC algorithm employed here follows directly from the theoretical result, being designed to approximate Gibbs-sampling in the theoretical, infinite-width tree structured Markov network. The degree of approximation is controlled by the value of a single hyperparameter that is also defined based on the theoretical result.
@@
The present paper stands apart from many other theoretical analyses of DNNs that view DNNs purely as function approximators and prove theorems about the quality of function approximation. Here we instead show that DNNs may be viewed as statistical models, specifically PGMs. This work is also different from the field of Bayesian neural networks, where the goal is to seek and model a probability distribution over neural network parameters. In our work, the neural network itself defines a joint probability distribution over its variables (nodes). Our work therefore is synergistic with Bayesian neural networks but more closely related to older work on learning stochastic neural networks via expectation maximization (EM) (Amari, 1995) or approximate EM (Song et al., 2016).
@@
Although the approach is different, our motivation is similar to that of Dutordoir et al. (2021) and Sun et al. (2020) in their work to link DNNs to deep Gaussian processes (GPs) (Damianou and Lawrence, 2013). By identifying the forward pass of a DNN with the mean of a deep GP layer, they aim to augment DNNs with advantages of GPs, notably the ability to quantify uncertainty over both output and latent nodes. What distinguishes our work from theirs is that we make the DNN-PGM approximation explicit and include all sigmoid DNNs, not just unsupervised belief networks or other specific cases.
All code needed to reproduce our experimental results may be found at https://github.com/engelhard-lab/DNN_TreePGM.
@@ 2 Background: Comparison to Bayesian Networks and Markov Networks
Syntactically a Bayesian network (BN) is a directed acyclic graph, like a neural network, whose nodes are random variables. Here we use capital letters to stand for random variables, and following Russell and Norvig (Russell and Norvig, 2020) and others, we take a statement written using such variables to be a claim for all specific settings of those variables. Semantically, a BN represents a full joint probability distribution over its variables as $P(\vec{V})=\prod_{i}P(V_{i}|pa(V_{i}))$, where $pa(V_{i})$ denotes the parents of variable $V_{i}$. If the conditional probability distributions (CPDs) $P(V_{i}|pa(V_{i}))$ are all logistic regression models, we refer to the network as a sigmoid BN.
@@
A DNN of any architecture is, like a Bayesian network, a directed acyclic graph. A sigmoid activation can be understood as a logistic model, thus giving a conditional probability distribution for a binary variable given its parents. Thus, there is a natural interpretation of a DNN with sigmoid activations as a Bayesian network (e.g., Bayesian belief network). Note, however, that when the DNN has multiple, stacked hidden nodes, the values calculated for those nodes in the DNN by its forward pass do not match the values of the corresponding hidden nodes in a Bayesian network. Instead, for the remainder of this paper, we adopt the view that the DNN’s forward pass might serve as an approximation to an underlying PGM and explore how said approximation can be precisely characterized. As reviewed in Appendix A, this Bayes net in turn is equivalent to (represents the same probability distribution) as a Markov network where every edge of weight $w$ from variable $A$ to variable $B$ has a potential of the following form:
@@ 3 The Construction of Tree-structured PGMs
Although both a binary pairwise Markov network (MN) and a Bayesian network (BN) share the same sigmoid functional structure as a DNN with sigmoid activations, it can be shown that the DNN does not in general define the same probability for the output variables given the input variables: forward propagation in the DNN is very fast but yields a different result than VE in the MN or BN, which can be much slower because the inference task is NP-complete. Therefore, if we take the distribution $\cal{D}$ defined by the BN or MN to be the correct meaning of the DNN, the DNN must be using an approximation $\cal{D^{\prime}}$ to $\cal{D}$. Procedurally, the approximation can be shown to be exactly the following: the DNN repeatedly treats the expectation of a variable $V$, given the values of $V$’s parents, as if it were the actual value of $V$. Thus previously binary variables in the Bayesian network view and binary features in the Markov network view become continuous. This procedural characterization of the approximation of $\cal{D^{\prime}}$ to $\cal{D}$ yields exactly the forward pass in the neural network within the space of a similarly structured PGM, yet, on its own, does not yield a precise joint distribution for said PGM. We instead prefer in the PGM literature to characterize approximate distributions such as $\cal{D^{\prime}}$ with an alternative PGM that precisely corresponds to $\cal{D^{\prime}}$; for example, in some variational methods we may remove edges from a PGM to obtain a simpler PGM in which inference is more efficient. Treewidth-1 (tree-structured or forest-structured) PGMs are among the most desirable because in those models, exact inference by VE or other algorithms becomes efficient. We seek to so characterize the DNN approximation here.
This approach aligns somewhat with the idea of the computation tree that has been used to explore the properties of belief propagation by expressing the relevant message passing operations in the form of a tree (Tatikonda and Jordan, 2002; Ihler et al., 2005; Weitz, 2006). Naturally the design of the tree structured PGM proposed here differs from the computation trees for belief propagation as we instead aim to capture the behavior of the forward pass of the neural network. Nonetheless, both methods share similar general approaches, the construction of a simpler approximate PGM, and aims, to better understand the theoretical behavior of an approximation to a separate original PGM.
@@
  3. c.
    Consider then the children of these nodes. Again, for each outgoing edge, make a copy of this node that is only connected to one child with that edge. In this step, for each copied node, we then also copy the entire subgraph formed by all ancestor nodes of the current node. Note that while weights across copies are tied, the values of the copies of any node are not tied. However, since we also copy the subtree of all input and intermediary hidden nodes relevant to a forward pass up to each copy, the probability of any of these copied nodes being true remains the same across copies (ignoring the influence of any information passed back from their children).
  4. d.
    We repeat this process across each layer until we have separate trees for each output node in the original deep neural network graph.
This process ultimately creates a graph whose undirected structure is a tree or forest. In the directed structure, trees converge at the output nodes. The probability of any copy of a latent node given the observed input (and ignoring any information passed back through a node’s descendant) is the same across all the copies, but when sampling, their values may not be.
@@
The preceding step alone is still not sufficient to accurately express the deep neural network as a PGM. Recall that in the probabilistic graphical model view of the approximation made by the DNN’s forward pass, the neural network effectively takes a local average, in place of its actual value, from the immediately previous nodes and passes that information only forward. The following additional step in the construction yields this same behavior. This next step of the construction creates $L$ copies of every non-output node in the network (starting at the output and moving backward) while also copying the entire ancestor subtrees of each of these nodes, as was done in step 1. The weight of a copied edges is then set to its original value divided by $L$. Note that this step results in a number of total copies that grows exponentially in the number of layers (i.e. $L$ copies in the $2$nd to last layer, $L^{2}$ copies in the layer before, etc). Detailed algorithms for the two steps in the construction of the infinite tree-structured PGM are presented in Appendix B. As $L$ approaches infinity, we show that both inference and the gradient in this PGM construction matches the forward pass and gradient in the neural network exactly.
@@
This second step in the construction can be thought of intuitively by considering the behavior of sampling in the Bayesian network view. Since we make $L$ copies of each node while also copying the subgraph of its ancestors, these copied nodes all share the same probabilities. As $L$ grows large, even if we sampled every copied node only once, we would expect the average value across these $L$ copies to match the probability of an individual copied node being true. Given that we set the new weights between these copies and their parents as the original weights divided by $L$, the sum of products (new weights times parent values) yields the average parent value multiplied by the original weight. As $L$ goes to infinity, we remove sampling bias and the result exactly matches the value of the sigmoid activation function of the neural network, where this expectation in the PGM view is passed repeatedly to the subsequent neurons. The formal proof of this result, based on variable elimination, is found in Appendix C. There, we show the following:
Theorem 3.1 (Matching Probabilities).
@@
In the PGM construction, as $L\to\infty$, $P(H=1|\vec{x})\to\sigma(\sum_{j=1}^{M}w_{j}g_{j}+\sum_{i}^{N}\theta_{i}\sigma(p_{i}))$, for an arbitrary latent node $H$ in the DNN that has observed parents $g_{1},...,g_{M}$ and latent parents $h_{1},...,h_{N}$ that are true with probabilities $\sigma(p_{1}),...,\sigma(p_{N})$. Here, $\sigma(\cdot)$ is the logistic sigmoid function and $w_{1},...,w_{M}$ and $\theta_{1},...,\theta_{N}$ are the weights on edges between these nodes and $H$.
The PGM of our construction is a Markov network that always has evidence at the nodes $\vec{X}$ corresponding to the input nodes of the neural network. As such, it is more specifically a conditional random field (CRF). Theorem 1 states the probability that a given node anywhere in the CRF is true given $\vec{X}$ equals the output of that same node in the neural network given input $\vec{X}$. The CRF may also have evidence at the nodes $\vec{Y}$ that correspond to the output nodes of the neural network. Given the form of its potentials, as illustrated at the end of Section 2, the features in the CRF’s loglinear form correspond exactly to the edges and are true if and only if the nodes on each end of the edge are true. It follows that the gradient of this CRF can be written as a vector with one entry for each feature $f$ corresponding to each weight $w$ of the neural network, of the form $P(f|\vec{X})-P(f|\vec{X},\vec{Y})$. Building on Theorem 1, this gradient of the CRF can be shown to be identical to the gradient of the cross-entropy loss in the neural network: the partial derivative of the cross-entropy loss with respect to the weight $w$ on an edge, or feature $f$ of the CRF, is $P(f|\vec{X})-P(f|\vec{X},\vec{Y})$. This result is more precisely stated below in Theorem 2 below, which is proven in Appendix D.
Theorem 3.2 (Matching Gradients).
In the PGM construction, as $L\to\infty$, the derivative of the marginal log-likelihood, where all hidden nodes have been summed out, with respect to a given weight exactly matches the derivative of the cross entropy loss in the neural network with respect to the equivalent weight in its structure.